OpenStudy (anonymous):

how do you put y=2(x+6)^2-8 in standard form?

OpenStudy (anonymous):

(x+6)^2= x^2+12x+36

OpenStudy (anonymous):

2x^2+24x+72

OpenStudy (anonymous):

-8= 2x^2+24x+64

OpenStudy (anonymous):

\[y=2x^2+24x+64\]

OpenStudy (anonymous):

Understand? By the way (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

its an identity

OpenStudy (anonymous):

@Bitmaximus thank you soooo much!